Additional Problem #2 Base Model Outcome and Evaluation

Cost Structure TC = 2.5Q2 + 10

Revenue Structure Q = 10 - .4P

Evaluation:

6. Quantity 7. Price 8. Consumer Surplus 9. Producer Surplus 10. Total Surplus 11. Dead-weight Loss
Answer #1

Inverse demand is P = 25 - 2.5Q

From the information provided, we have Marginal cost MC = 5Q and marginal revenue MR = 25 - 5Q

At profit maximizing output MR = MC

25 - 5Q = 5Q

Q = 25/10 = 2.5 units

Price = 25 - 2.5*2.5 = $18.75

Consumer surplus = 0.5*(Max price - price paid)*qty = 0.5*(25 - 18.75)*2.5 = 7.8125

Producer surplus = area of trapezium so formed = 0.5*(18.75 - 0 + 18.75 - 12.5)*2.5 = 31.25

Total surplus = CS + PS = 39.0625

Dead weight loss = 0.5*(price - MC)*qty = 0.5*(18.75 - 12.5)*(3.33 - 2.5) = 2.60